Calculate the activity of K40 in 100kg man assuming that 0.35% of the body weight is potassium. The abundance of K40 is 0.012%, its half life is 1.31 109 years [ Avagadro number, N= 6.023 106] .
1
Expert's answer
2019-04-08T12:56:48-0400

The number of atoms 40K N = 6*1023*350*12*10-5/40 = 6.3*1020, then the activity A = λN = 0.69*6.3*1020/(1.31*109*365*24*3.6*103) = 1.05*104 (s-1).